OpenStudy (anonymous):

megan is 5 years older than sandy, if 4 years ago megan was 3 times sandys age how old is sandy?

OpenStudy (anonymous):

PRESENT AGE OF SANDY BE X AND THAT OF MEGAN BE X+5 4 YRS AGO AGE OF SANDY = X-4 AND THAT OF MEGAN IS X+5-4=X+1 GIVEN X+1=3(X-4) and so on...